OpenStudy (anonymous):

Find the zeros of the function: f(x) = x2 - 5x - 6. Enter the smallest zero value.

OpenStudy (ash2326):

let's factor f(x) \[ f(x)=x^2-5x-6\] \[ f(x)= x^2-6x+x-6\] \[f(x)= x(x-6)+1(x-6)\] \[f(x)= (x+1)(x-6)\] so the roots are \[x=-1, 6\] now airricuhh you can find the smaller one, can't you?

OpenStudy (anonymous):

thank you ash <3

OpenStudy (ash2326):

welcome airricuhh but you should try now yourself these type of examples, if you get stuck , do tell me:)